Mike and hassan together harvested 19 mangosteens.if Hassan has 3 less of the fruit than mike, find the number of mangosteens mike has

Answer by amarjeeth123(569)

Let the number of mangosteens of Mike be x.
Then the number of mangosteens of Hassan is (x-3).
x+(x-3)=19
2x-3=19
2x=19+3
2x=22
x=11
Mike has 11 mangosteens.